OpenStudy (anonymous):

what is the graph of -3x-8y=24?

OpenStudy (jdoe0001):

solve for "y", what do you get?

OpenStudy (anonymous):

straight line graph rearrange in the form of y=mx+C where m is gradient and C is intercept -3x-8y = 24 -8y=24+3x y= 3x/-8 + 24/-8 y=-3x/8 -3 Slope -3/8 intercept -3

OpenStudy (anonymous):

@jdoe0001 okay so i solved for x= -8 and y=3

OpenStudy (anonymous):

nvm thank you both @jdoe0001 @anette

OpenStudy (charlotte123):

u still need help? :)

OpenStudy (anonymous):

how do you put this in on a calculator though? @charlotte123

OpenStudy (charlotte123):

I don't think you can do that :P I use wolframalpha for problems like these, especially when u don't have a graph available: http://www.wolframalpha.com/input/?i=-3x-8y%3D24 TA DA

OpenStudy (anonymous):

okay thanks @charlotte123

OpenStudy (anonymous):

welcome

OpenStudy (charlotte123):

no prob! :)